OpenStudy (laviajmani):

Solve the following equation for x e^(3x^2 - 14x) = e^5

OpenStudy (styxer):

You can cancel out the "e" and just calculate (3x^2-14x)=5 , it will result in bhaskara 3x^2-14x-5=0 x can be either -1/3 or 5

OpenStudy (sweetburger):

Yes considering both sides of the equation have the same base you can say \[3x^2-14x=5\]